The longest wavelength of light with enough energy to break the Br–Br bond in Br2g) is 620 nm.

Calculate the frequency of the light.
Answer #1

The relationship between wavelength(\lambda) of the light and frequency (\nu) of the light is given by   \nu = C/\lambda where C is velocity of light

= 3.0*10^8/620*10^-9

= 4.84*10^14 Hzs.

​​​​​​​Ans: frequency of light = 4.84*10^14 Hz